OpenStudy (anonymous):

someone PLEASE HELP ME!!!!! I REALLY NEED HELP!! Solve the equation for y: -36x + 6y = 18. y = 3 + 6x y = 3 – 6x y = 3 + 36x y = 12 + 36x

OpenStudy (anonymous):

THANK YOU SOOOOOOO MUCH!!!!!!!!!!!!!

OpenStudy (anonymous):

\[6y=18(2x+1)\] \[y=3(2x+1)\] \[y=6x+3\]

OpenStudy (anonymous):

-36x + 6y = 18 Let's move the -36x to the other side to isolate it. 6y = 36x + 18 Now lets divide by 6 on all terms on the RHS. y = 36x/6 + 3 y = 6x + 3 It helps to see how it's done, and not just get a final answer. This is how you learn :)

OpenStudy (anonymous):

THANKS!!!!!!!!!!!!!!!!!